OpenStudy (prince1342):

Quadrilateral OPQR is inscribed in circle N, as shown below. Which of the following could be used to calculate the measure of ∠QRO?

OpenStudy (prince1342):

m∠QPO + (x + 16)° + (6x − 4)° = 360° m∠QPO = (x + 16)° + (6x − 4)° m∠QPO + (2x + 16)° = 180° m∠QPO = (6x − 4)° + (2x + 16)°

OpenStudy (error1603):

I'll help!

OpenStudy (prince1342):

Uh, thanks!

OpenStudy (error1603):

Why Uh?

OpenStudy (prince1342):

Because of the picture of the old people lol

OpenStudy (prince1342):

Can you help please? @mathmale

OpenStudy (prince1342):

@ganeshie8